24-carat gold is pure gold.
18 carat gold is 3/4 gold.
20- carat gold is 5/8 gold.
the ratio of pure gold in 18 carat gold to pure gold in 20 carat gold is ?

Found 2 solutions by noahhamstra, abdullahkhawer:
Answer by noahhamstra(1)
Answer by abdullahkhawer(5)

Question:
The ratio of pure gold in 18 carat gold to pure gold in 20 carat gold is 3/4 : 5/8 = (3/4) * (8/5) because whenever a value say 'a' is being divided by some value say 'b' then 'a' can be multiplied by the reciprocal of 'b'.
Therefore, (3/4) * (8/5) = 24/20 = 6/5
--------
Answer:
The ratio of pure gold in 18 carat gold to pure gold in 20 carat gold is 6/5.
--------
Remarks:
I hope that helped you. :)